OpenStudy (anonymous):

how do i find tan sqrt3/3?

OpenStudy (anonymous):

\[\tan \sqrt{3}/3\]

OpenStudy (anonymous):

sorry, tan (theta) sqrt3/3

OpenStudy (anonymous):

tan theta = sqrt 3 /3 ?

OpenStudy (anonymous):

yeah, there are aparently 6 solutions

OpenStudy (anonymous):

like a formula solution

OpenStudy (anonymous):

how to you find theta you mean? use arctan aka tan^-1

OpenStudy (anonymous):

well 5pi/6 is one of the solutions so how would i write a formula solution?

OpenStudy (anonymous):

sqrt3/3 = sqrt3/(sqrt3)^2 = 1/sqrt3 tan theta=1/sqrt3 so the answer is 30

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

(1/6)*Pi and there ar an infinite amnt of solotions. Cuzz it is a trig function.

OpenStudy (anonymous):

other wise just use calc in radian mode to get tan(3√/3)=0.6513878868

OpenStudy (anonymous):

well thats not what im looking for, thank you. the answer on the back of my book says theta=5pi/6+kpi and i dont know how they got that answer

OpenStudy (anonymous):

the + kpi just means that every k_pi it repeats.

OpenStudy (anonymous):

how do you get 5pi/6?

OpenStudy (anonymous):

well when I and edwardp did it we used calculators to find arctan and it gave us (1/6)*Pi which is pi/6 Put since it is repeating. I assume that 5pi/6 is also an answer. Just my guess.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah